# Few-Shot Adversarial Low-Rank Fine-Tuning of Vision-Language Models

### v2

@@ 5 Empirical Results
Table 1: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold.
@@
1 C-AVP [37] 43.62 17.94 46.60 11.07 85.73 50.33 26.97 12.93 24.43 5.23 57.60 22.73 63.10 29.70 3.37 0.40 41.20 11.10
APT [24] 59.07 5.08 49.30 1.30 84.77 26.90 41.67 3.83 56.57 0.83 70.23 0.60 61.97 2.10 54.50 3.87 53.53 1.23
AdvPT [25] 29.96 1.44 20.17 0.43 62.97 7.60 16.73 2.60 13.27 0.00 37.93 0.13 33.97 0.43 27.03 0.00 27.57 0.37
AdvMaPLe [30] 33.52 11.38 49.27 14.60 85.53 48.37 13.63 2.93 5.27 0.30 30.67 4.97 1.40 0.10 32.70 7.07 49.70 12.67
AdvVLP [34] 32.82 11.78 49.00 15.53 85.43 48.47 15.97 4.77 1.07 0.77 29.63 3.83 19.77 6.57 11.83 1.73 49.83 12.60
FAP [34] 40.14 10.22 49.90 15.40 83.53 41.13 18.40 2.40 31.67 1.43 49.23 3.47 10.40 0.53 28.50 2.43 49.53 14.93
AdvCLIP-LoRA1 67.18 23.9 50.86 15.89 90.79 59.31 47.70 14.24 74.22 11.81 85.80 29.14 64.27 18.64 63.73 17.31 60.05 24.90
AdvCLIP-LoRA2 66.12 24.24 51.94 17.19 87.42 58.22 48.35 15.31 73.34 11.94 83.18 29.30 63.74 19.04 61.70 16.92 59.30 25.99
AdvCLIP-LoRA5 64.22 26.32 50.29 17.53 85.56 60.93 47.75 16.96 72.82 15.01 79.39 32.92 60.94 19.57 58.90 19.83 58.10 27.78
AdvCLIP-LoRA10 61.70 27.12 47.57 17.66 83.61 60.41 47.22 16.43 69.66 18.10 74.93 32.71 60.13 22.05 53.85 21.02 56.62 28.61
AdvCLIP-LoRA25 53.32 26.28 36.96 14.74 79.07 61.46 45.74 18.79 55.32 17.59 62.77 30.36 49.86 20.30 45.33 19.67 51.52 27.29
AdvCLIP-LoRA50 34.78 19.41 27.66 10.44 68.15 57.69 39.83 18.91 17.17 4.47 29.38 14.80 15.14 8.57 36.06 17.71 44.88 22.71
AdvCLIP-LoRA75 27.75 16.53 22.18 8.26 64.54 54.52 36.23 20.15 10.20 2.67 17.66 9.27 3.74 2.80 29.08 14.17 38.39 20.43
AdvCLIP-LoRA100 23.23 14.81 15.51 5.57 62.27 52.86 30.61 19.50 8.15 2.17 10.00 6.57 3.21 3.13 25.27 12.61 30.80 16.08
4 C-AVP [37] 43.10 16.40 49.80 11.13 90.17 52.50 18.77 9.27 22.73 4.57 57.80 16.20 55.97 23.73 1.07 0.80 48.47 13.03
APT [24] 66.37 6.04 50.90 1.40 90.77 26.67 51.33 6.33 54.80 1.63 71.83 2.10 82.40 4.23 66.53 3.03 62.37 2.90
AdvPT [25] 35.32 2.07 23.40 1.33 64.97 7.30 31.70 4.37 15.23 0.37 44.13 1.73 41.97 0.63 31.17 0.47 29.97 0.40
AdvMaPLe [30] 51.01 21.61 51.27 19.00 89.53 59.40 6.43 2.40 60.00 14.83 30.70 9.03 52.20 25.37 59.73 21.30 58.23 21.53
AdvVLP [34] 55.18 23.40 51.30 19.37 89.37 59.07 22.97 10.33 41.50 11.20 67.43 18.47 51.00 25.80 59.97 21.77 57.90 21.17
FAP [34] 57.51 24.6 51.53 19.60 87.57 57.33 31.27 8.07 59.37 18.37 42.10 9.30 73.13 38.77 58.50 22.13 56.60 23.20
AdvCLIP-LoRA1 76.11 31.89 62.53 21.89 93.59 69.41 58.69 21.87 77.21 15.79 86.86 35.65 86.20 35.16 74.20 24.11 69.64 31.26
AdvCLIP-LoRA2 76.24 34.05 62.47 22.91 93.35 71.40 59.69 23.88 77.29 18.10 87.68 37.48 85.91 39.26 73.91 26.51 69.60 32.84
AdvCLIP-LoRA5 75.75 37.12 62.11 24.69 93.10 73.59 57.51 27.13 76.49 20.91 88.12 40.58 85.79 45.11 73.80 29.58 69.09 35.39
AdvCLIP-LoRA10 75.09 38.07 61.41 25.43 93.27 74.20 57.57 27.54 75.11 23.01 87.54 41.40 85.51 48.40 71.82 28.26 68.49 36.29
AdvCLIP-LoRA25 72.93 39.26 58.08 25.56 93.18 75.90 53.90 27.84 73.33 26.25 87.11 41.26 81.24 50.51 70.53 30.82 66.03 36.94
AdvCLIP-LoRA35 71.52 39.12 56.40 24.58 93.10 78.30 53.07 26.89 71.65 25.47 85.88 41.16 79.54 48.84 67.88 31.03 64.66 36.68
AdvCLIP-LoRA50 69.21 38.42 54.52 23.33 92.41 76.84 51.65 26.65 68.55 24.57 83.95 39.98 75.23 48.40 64.34 31.96 63.03 35.65
AdvCLIP-LoRA75 63.54 35.21 46.98 19.83 91.20 75.70 48.17 26.71 57.67 19.82 82.77 37.53 64.35 40.48 57.71 28.36 59.46 33.23
AdvCLIP-LoRA100 52.08 30.12 20.24 10.42 88.64 73.59 44.98 24.70 22.08 7.55 81.22 35.57 53.35 33.37 53.56 26.43 52.54 29.33
16 C-AVP [37] 41.90 17.13 46.27 12.77 90.40 52.60 29.20 13.87 1.07 0.80 56.40 16.43 56.17 22.03 0.97 0.93 54.70 17.63
APT [24] 71.05 8.35 52.63 2.07 92.93 30.23 54.93 10.47 62.50 2.63 83.70 4.40 86.63 8.97 69.40 4.40 65.67 3.67
AdvPT [25] 40.94 2.68 24.53 1.47 68.70 9.63 43.77 5.70 18.47 0.73 46.27 0.23 56.03 0.80 36.60 0.53 33.13 2.37
AdvMaPLe [30] 71.48 38.11 52.93 21.90 92.17 68.63 57.93 32.17 65.13 25.27 83.27 36.87 87.87 58.70 68.97 31.67 63.57 29.70
AdvVLP [34] 68.76 37.01 53.23 22.10 92.37 67.97 57.53 32.73 43.30 16.50 82.93 35.57 87.70 58.70 69.10 32.80 63.90 29.70
FAP [34] 69.88 39.22 52.53 22.90 91.10 67.33 55.17 31.33 64.03 26.67 81.90 41.00 86.27 61.47 65.70 32.80 62.37 30.27
AdvCLIP-LoRA1 80.91 35.09 66.83 24.88 95.29 72.66 66.55 25.30 78.68 17.07 88.74 33.17 95.62 48.56 81.68 25.17 73.87 33.88
AdvCLIP-LoRA2 80.96 38.25 66.74 26.03 95.25 75.38 66.08 29.61 78.52 19.71 89.21 36.55 95.82 52.82 82.10 28.60 73.99 37.26
AdvCLIP-LoRA5 80.61 40.24 66.59 28.11 95.74 76.67 65.37 30.20 78.36 22.14 89.13 37.56 95.86 56.03 79.99 32.09 73.84 39.15
AdvCLIP-LoRA10 80.12 41.53 66.06 29.48 95.38 76.80 64.60 31.74 77.57 24.21 88.66 37.53 95.53 57.53 79.70 34.58 73.49 40.35
AdvCLIP-LoRA25 78.73 42.40 64.52 30.50 95.42 78.26 63.00 31.86 75.34 24.50 86.86 38.38 94.32 58.95 77.82 34.63 72.55 42.14
AdvCLIP-LoRA35 77.49 41.97 63.06 29.81 95.21 78.78 61.88 30.97 73.74 23.61 86.07 37.88 93.10 59.11 75.15 33.70 71.70 41.89
AdvCLIP-LoRA50 75.74 39.24 61.45 28.74 94.81 78.95 58.87 29.91 70.75 20.81 85.01 35.60 92.12 50.99 72.61 27.73 70.27 40.17
AdvCLIP-LoRA75 72.41 35.04 58.38 25.86 93.55 74.73 55.08 26.77 66.56 18.21 82.15 29.93 89.69 47.06 69.36 23.10 64.54 34.63
AdvCLIP-LoRA100 68.39 31.50 46.58 21.85 91.72 71.60 53.43 25.06 61.82 13.60 80.19 25.29 86.85 40.32 65.13 22.10 61.40 31.17
5.1 Experiments Setup
Practical Implementation Notes for AdvCLIP-LoRA. We highlight a few minor differences between the theoretically analyzed version of AdvCLIP-LoRA and the implementation used in our experimental section. Such discrepancies are common in the literature, reflecting the gap between theoretical analysis and practical scenarios. 1) The proposed algorithm is guaranteed to visit an $\epsilon$-stationary point within a specified number of iterations by returning $\hat{A}$ and $\hat{B}$, sampled uniformly from the set $\{(A_{t},B_{t})\}_{t=1}^{T}$. However, this does not imply that the final iterate or the model achieving the best validation accuracy corresponds to the $\epsilon$-stationary point. In our experiments, we report the test accuracy of the model that achieves the highest accuracy on the validation set. 2) At the beginning of each iteration in AdvCLIP-LoRA i.i.d. samples are drawn randomly for the dataset. However, in our implementation, we fix a minibatch and use the entire training set in every iteration. 3) As discussed in Section 3.2, AdvCLIP-LoRA updates $A$, $B$, and $\delta$ once per iteration, and we have shown that this procedure guarantees convergence to a stationary solution. However, our empirical findings suggest that performing multiple updates and projections on $\delta$ per iteration can enhance robustness, though it may lead to a slight reduction in clean accuracy on some datasets. This gain in robustness stems from a more precise approximation of the inner maximization in the minimax formulation, allowing the model to better anticipate adversarial perturbations. We denote the number of $\delta$-updates per iteration by $\tau$, and conduct extensive experiments to evaluate the impact of varying $\tau$.
Datasets. To evaluate the proposed method, we follow prior works [27, 26] and utilize a diverse set of 8 image recognition datasets spanning multiple vision tasks. The datasets include two generic object recognition datasets: ImageNet-1K [8] and Caltech101 [38]; a texture recognition dataset: DTD [39]; four fine-grained object recognition datasets: OxfordPets [40], Flowers102 [41], and Food101 [42]; a scene recognition dataset: SUN397 [43]; and an action recognition dataset: UCF101 [44].
Implementation Details. To evaluate the performance of our proposed method, we conduct extensive experiments comparing AdvCLIP-LoRA with CLIP-LoRA, using ViT-B/16 and ViT-B/32 backbones. Our experimental setup closely follows that of [16] to ensure a fair comparison. LoRA is applied to both the vision and text encoders of CLIP, with the rank of $2$ and a dropout layer with $p=0.25$. We use a batch size of $16$ for ImageNet-1K and $32$ for the other datasets. The number of training iterations is set to $500\times N/K$. All experiments are conducted on NVIDIA A6000 and V100 GPUs.
Learning Rates. For the low-rank matrices, we use a learning rate of $2*10^{-4}$ with a cosine decay scheduler. Selecting an appropriate learning rate for $\delta$ posed a challenge, since the gradients for $\delta$ were too small to enable effective updates at the beginning of the training. To address this, we adopt a larger, adaptive learning rate defined as $\eta_{\delta}=0.05/\|\delta_{t}\|_{2}$, which scales inversely with the magnitude of $\delta_{t}$. This adjustment amplifies updates, improving early-stage learning. It can also be viewed as an implicit data augmentation strategy, introducing noise into the learning process. The learning rate $\eta_{\delta}$ then dynamically decays during training, eventually stabilizing at $0.05$ after $300$ iterations.
Adversarial Attacks. For adversarial training, we define the projection set for updating $\delta$ as an $\ell_{\infty}$-ball with a radius of $\epsilon\in\{1/255,10/255\}$ across all datasets. To evaluate adversarial robustness, we implement two standard attack methods: FGSM [21] and PGD [23]. For FGSM, we set $\epsilon=10/255$, while for PGD, we use $\epsilon\in\{1/255,2/255\}$ with a total of 20 attack iterations. Following the setup in [23], the step size for PGD is set to $\alpha=\frac{2.5\cdot\epsilon}{\texttt{Number of Iterations}}$.
5.2 Comparative Analysis of AdvCLIP-LoRA with Robust PEFT Methods
We conduct a thorough comparison between AdvCLIP-LoRA and competitive baselines using ViT-B/32 as the backbone across 8 datasets. Table 1 presents the results for 1-, 4-, and 16-shot settings. Complete results and corresponding plots are provided in Table 3 and Fig. B in the Appendix. Our baselines include state-of-the-art methods such as C-AVP [1], APT [2], AdvPT [3], AdvMaPLe [4], AdvVLP [5], and FAP [5]. We follow the setup in FAP [5] for measuring adversarial robustness using a 2-step PGD attack with an $\ell_{\infty}$ norm, a perturbation bound of $\epsilon=1/255$, and a step size of $\alpha=1/255$. For a fair comparison, we report results for $\tau=1,2$ (matching the baselines) and also explore higher $\tau$ values to illustrate the enhanced adversarial fine-tuning capabilities of our method.
On average, over all datasets:
  • Clean Accuracy: AdvCLIP-LoRA with $\tau=2$ outperforms the best baseline (FAP) by 25.98, 16.66, 18.75, 11.35, and 11.08 percentage points for 1, 2, 4, 8, and 16 shots, respectively.
  • Robust Accuracy: AdvCLIP-LoRA with $\tau=2$ exceeds FAP by 14.02, 11.71, 9.45, and 0.87 percentage points for 1, 2, 4, and 8 shots; for 16 shots, it is only 0.97 percentage points below FAP.
We also note that in our method, increasing $\tau$ up to a moderate threshold (about 25 ) can further improve the clean/robust trade-off for most tasks, whereas the effect on other baselines is unclear and would require more extensive retuning.
5.3 Practical Guidance on Selecting the Hyperparameter $\tau$
According to Figure 1 (full version in Fig 3 in the Appendix), our observations on the choice of $\tau$ are as follows:
  • Optimal Range for Robustness: For most tasks, robust accuracy continues to improve up to $\tau=25$ for few-shot settings ( $2,4,8$, and 16 shots) and up to $\tau=10$ for the 1-shot setting.
  • Point of Diminishing Returns: Beyond $\tau=25$, we observe that for the majority of datasets, both clean and robust accuracy begin to decrease. This suggests the model may start to overfit to the specific adversarial examples generated during training, harming its generalizability.
  • Dataset-Specific Behavior: We note some tasks, like Caltech101, can benefit from even larger $\tau$ values (up to 35 ), highlighting that the optimal $\tau$ can have some data dependency.
  • Clear Trade-off: The results confirm a clear trade-off where larger $\tau$ values consistently improve robustness up to a point, but at a predictable cost to clean accuracy.
From this analysis, we recommend setting $\tau$ between 10 and 25 as a practical choice for achieving a strong balance between clean and robust accuracy in few-shot learning.

### v3

@@ B.1 Adversarial Few-Shot Learning
Table 5: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone. Top-1 accuracy averaged over 3 random seeds is reported.
@@
16 AdvVP Mao et al. 41.90 17.13 46.27 12.77 90.40 52.60 29.20 13.87 1.07 0.80 56.40 16.43 56.17 22.03 0.97 0.93 54.70 17.63
APT Li et al. (2024) 71.05 8.35 52.63 2.07 92.93 30.23 54.93 10.47 62.50 2.63 83.70 4.40 86.63 8.97 69.40 4.40 65.67 3.67
AdvPT Zhang et al. (2024) 40.94 2.68 24.53 1.47 68.70 9.63 43.77 5.70 18.47 0.73 46.27 0.23 56.03 0.80 36.60 0.53 33.13 2.37
AdvMaPLe Khattak et al. (2023) 71.48 38.11 52.93 21.90 92.17 68.63 57.93 32.17 65.13 25.27 83.27 36.87 87.87 58.70 68.97 31.67 63.57 29.70
AdvVLP Zhou et al. (2024) 68.76 37.01 53.23 22.10 92.37 67.97 57.53 32.73 43.30 16.50 82.93 35.57 87.70 58.70 69.10 32.80 63.90 29.70
FAP Zhou et al. (2024) 69.88 39.22 52.53 22.90 91.10 67.33 55.17 31.33 64.03 26.67 81.90 41.00 86.27 61.47 65.70 32.80 62.37 30.27
AdvCLIP-LoRA ($\tau=1$) 81.26 36.08 68.42 25.05 95.29 72.66 67.49 26.95 77.88 16.83 88.25 32.52 96.67 52.46 81.89 30.00 74.17 32.20
AdvCLIP-LoRA ($\tau=2$) 81.23 37.28 68.38 25.86 94.93 72.98 67.67 28.37 77.81 17.76 88.44 34.29 96.47 54.69 81.87 30.74 74.23 33.52
AdvCLIP-LoRA ($\tau=5$) 81.02 38.63 68.28 27.05 95.05 75.42 67.20 28.31 77.65 18.98 88.01 34.59 96.39 55.62 81.34 34.10 74.22 34.98
AdvCLIP-LoRA ($\tau=10$) 81.04 39.53 68.12 28.06 94.97 76.63 67.20 29.96 77.32 19.55 88.20 36.33 96.22 55.83 82.18 34.10 74.11 35.76
AdvCLIP-LoRA ($\tau=15$) 80.95 40.68 67.97 28.59 94.93 76.96 67.32 31.97 77.38 20.40 87.68 37.37 96.31 57.49 82.10 36.16 73.91 36.51
AdvCLIP-LoRA ($\tau=20$) 80.79 40.73 67.90 28.85 94.97 77.00 66.84 31.74 77.22 20.37 87.87 37.61 96.22 57.21 81.26 36.08 74.03 37.01
AdvCLIP-LoRA ($\tau=25$) 80.83 41.06 67.88 29.03 94.93 76.96 67.32 32.39 76.97 20.82 87.44 37.78 96.35 57.33 81.68 36.64 74.09 37.55
AdvCLIP-LoRA ($\tau=30$) 80.70 40.93 67.82 29.25 95.21 77.04 66.73 30.73 76.86 20.72 87.54 38.05 96.31 58.18 81.15 36.08 73.98 37.42
Relative Improvement  +$13.37$  +$0.79$  +$27.97$  +$22.53$  +$2.2$  +$11.66$  +$16$ 8.46-8.46  +$18.72$ 26.7-26.7  +$5.38$ 11.39-11.39  +$9.5$ 9.18-9.18  +$18.41$  +$3.96$  +$12.85$  +$18.14$
8 AdvVP Mao et al. 43.24 17.21 46.37 11.90 91.20 50.33 23.63 11.47 1.00 0.83 57.43 17.33 55.50 23.57 18.27 4.93 52.53 17.30
APT Li et al. (2024) 69.76 7.56 52.03 1.80 92.37 30.83 54.43 8.70 61.57 2.33 82.87 3.10 84.00 6.00 66.53 4.30 64.30 3.40
AdvPT Zhang et al. (2024) 38.50 2.39 24.30 0.87 68.07 10.10 37.47 4.20 16.97 0.10 44.20 0.27 51.13 0.87 33.43 1.40 32.47 1.33
AdvMaPLe Khattak et al. (2023) 62.90 30.45 52.13 20.60 90.63 63.80 33.20 16.97 62.70 20.13 55.60 21.07 83.10 48.80 64.33 25.93 61.50 26.30
AdvVLP Zhou et al. (2024) 68.32 32.87 52.83 20.97 90.17 63.13 51.83 25.77 61.73 19.33 80.67 29.63 83.90 50.90 64.07 26.97 61.33 26.23
FAP Zhou et al. (2024) 67.23 34.26 52.17 21.53 89.63 62.50 52.13 25.77 61.80 23.20 79.47 34.57 81.53 52.63 60.70 26.67 60.40 27.23
AdvCLIP-LoRA ($\tau=1$) 79.56 34.36 67.24 23.65 94.28 70.75 64.72 25.12 77.17 15.79 87.95 32.57 92.73 48.52 80.12 27.70 72.26 30.81
AdvCLIP-LoRA ($\tau=2$) 79.32 35.44 67.11 24.49 94.60 72.09 63.24 26.42 77.03 16.93 87.71 33.63 92.49 48.68 80.31 29.16 72.09 32.08
AdvCLIP-LoRA ($\tau=5$) 79.27 36.73 67.16 25.68 94.56 72.90 63.36 27.90 76.77 18.76 87.54 34.83 92.61 49.98 80.20 30.69 71.97 33.09
AdvCLIP-LoRA ($\tau=10$) 79.19 37.93 67.00 26.61 94.20 74.24 63.06 28.90 76.32 19.87 87.76 35.35 92.57 51.48 80.41 32.59 72.17 34.38
AdvCLIP-LoRA ($\tau=15$) 79.02 38.33 66.74 27.08 94.28 74.48 63.00 29.43 76.13 20.33 87.41 36.09 92.53 51.81 80.23 32.78 71.87 34.62
AdvCLIP-LoRA ($\tau=20$) 78.91 38.54 66.67 27.36 93.87 75.38 63.12 29.67 75.78 20.57 87.08 35.38 92.61 52.42 80.02 32.33 72.12 35.22
AdvCLIP-LoRA ($\tau=25$) 78.80 38.88 66.64 27.61 94.00 75.09 62.59 30.56 75.91 21.31 87.00 35.40 92.33 52.09 79.83 33.31 72.11 35.68
AdvCLIP-LoRA ($\tau=30$) 78.84 39.05 66.61 27.82 93.91 75.33 62.94 30.44 75.71 21.33 87.27 36.33 92.12 52.42 80.07 33.10 72.11 35.64
Relative Improvement  +$13.52$  +$10.71$  +$26.82$  +$23.59$  +$1.98$  +$16.36$  +$15.86$  +$12.15$  +$21.72$ 14.35-14.35  +$5.9$  +$2.26$  +$10.2$ 2.19-2.19  +$20.86$  +$20.84$  +$12.24$  +$26.26$
4 AdvVP Mao et al. 43.10 16.40 49.80 11.13 90.17 52.50 18.77 9.27 22.73 4.57 57.80 16.20 55.97 23.73 1.07 0.80 48.47 13.03
APT Li et al. (2024) 66.37 6.04 50.90 1.40 90.77 26.67 51.33 6.33 54.80 1.63 71.83 2.10 82.40 4.23 66.53 3.03 62.37 2.90
AdvPT Zhang et al. (2024) 35.32 2.07 23.40 1.33 64.97 7.30 31.70 4.37 15.23 0.37 44.13 1.73 41.97 0.63 31.17 0.47 29.97 0.40
AdvMaPLe Khattak et al. (2023) 51.01 21.61 51.27 19.00 89.53 59.40 6.43 2.40 60.00 14.83 30.70 9.03 52.20 25.37 59.73 21.30 58.23 21.53
AdvVLP Zhou et al. (2024) 55.18 23.40 51.30 19.37 89.37 59.07 22.97 10.33 41.50 11.20 67.43 18.47 51.00 25.80 59.97 21.77 57.90 21.17
FAP Zhou et al. (2024) 57.51 24.60 51.53 19.60 87.57 57.33 31.27 8.07 59.37 23.20 79.47 34.57 81.53 52.63 60.70 26.67 60.40 27.23
AdvCLIP-LoRA ($\tau=1$) 77.88 33.70 66.38 22.92 94.08 69.78 61.17 26.36 75.91 16.40 87.05 32.22 90.99 48.27 76.37 24.08 71.05 29.55
AdvCLIP-LoRA ($\tau=2$) 77.94 34.74 66.34 23.78 93.96 71.03 62.41 26.36 75.80 17.69 87.03 32.98 90.70 48.72 76.18 26.22 71.09 31.11
AdvCLIP-LoRA ($\tau=5$) 77.71 36.01 66.10 24.84 93.87 72.21 61.70 27.96 75.41 18.76 86.78 33.91 90.58 50.95 76.13 26.86 71.08 32.62
AdvCLIP-LoRA ($\tau=10$) 77.57 37.36 65.96 25.58 93.91 73.59 61.11 28.43 75.06 20.48 87.00 35.90 90.17 52.33 76.37 28.97 71.00 33.57
AdvCLIP-LoRA ($\tau=15$) 77.41 37.80 65.91 26.15 94.20 73.91 61.11 29.14 74.82 20.99 86.97 36.20 89.93 52.54 75.55 29.58 70.82 33.85
AdvCLIP-LoRA ($\tau=20$) 77.20 38.02 65.87 26.50 93.75 73.75 60.28 29.14 74.49 21.13 86.75 35.30 90.01 53.88 75.65 30.06 70.81 34.43
AdvCLIP-LoRA ($\tau=25$) 77.10 38.19 65.81 26.67 93.91 74.32 60.34 29.20 74.62 21.83 85.75 35.40 90.05 53.02 75.44 30.61 70.86 34.49
AdvCLIP-LoRA ($\tau=30$) 77.16 38.59 65.77 26.88 93.59 74.77 60.22 29.61 74.42 21.86 86.40 36.47 90.38 53.55 75.89 30.69 70.64 34.90
Relative Improvement  +$16.88$  +$51.87$  +$28$  +$30.51$  +$3.46$  +$23.89$  +$19.05$  +$175.22$  +$26.43$ 11.72-11.72  +$9.48$  +$3.85$  +$9.43$ 0.57-0.57  +$14.79$  +$8.62$  +$13.84$  +$23.28$
2 AdvVP Mao et al. 39.24 16.23 46.23 10.90 91.25 55.23 14.27 6.93 1.05 0.10 47.13 15.10 61.47 26.93 1.73 1.07 50.77 13.57
APT Li et al. (2024) 63.56 5.90 48.83 1.03 89.70 31.70 45.57 4.27 60.17 0.87 72.87 1.07 67.17 3.10 65.00 3.10 59.20 2.03
AdvPT Zhang et al. (2024) 32.47 1.76 22.37 0.77 66.07 8.33 24.57 2.43 11.13 0.17 39.17 1.17 38.47 0.23 29.37 0.23 28.57 0.77
AdvMaPLe Khattak et al. (2023) 39.09 15.58 49.97 17.13 88.00 56.20 16.53 4.20 3.10 0.67 34.03 6.87 46.17 17.00 21.17 6.20 53.73 16.33
AdvVLP Zhou et al. (2024) 42.79 17.76 50.53 17.50 87.60 55.33 18.33 7.17 1.53 1.10 31.27 7.07 62.43 25.17 36.83 11.43 53.77 17.33
FAP Zhou et al. (2024) 55.18 18.14 48.53 17.83 87.73 53.90 18.40 4.33 56.90 10.53 64.23 12.67 53.10 19.57 58.50 7.03 54.07 19.30
AdvCLIP-LoRA ($\tau=1$) 75.89 32.04 65.69 21.70 93.55 69.41 57.98 23.88 76.04 15.74 86.18 33.99 84.86 40.19 74.52 23.50 68.29 27.90
AdvCLIP-LoRA ($\tau=2$) 75.75 32.55 65.74 22.58 93.35 69.94 57.15 23.52 76.00 16.92 86.32 34.31 84.57 39.14 74.39 24.95 68.50 29.01
AdvCLIP-LoRA ($\tau=5$) 75.73 34.27 65.60 24.03 93.27 71.76 57.51 24.47 75.64 18.46 86.43 35.95 85.18 41.78 73.57 26.51 68.67 31.20
AdvCLIP-LoRA ($\tau=10$) 75.58 35.35 65.47 24.80 93.23 73.47 57.45 25.24 75.22 19.48 86.40 36.79 85.02 43.65 73.33 27.17 68.49 32.24
AdvCLIP-LoRA ($\tau=15$) 75.70 35.94 65.36 25.37 93.39 73.51 58.22 26.30 75.07 20.26 86.02 37.20 85.51 44.09 73.51 27.91 68.56 32.90
AdvCLIP-LoRA ($\tau=20$) 75.56 36.29 65.30 25.59 93.51 73.96 57.39 26.36 74.91 20.56 86.43 37.72 85.10 44.86 73.28 28.05 68.59 33.24
AdvCLIP-LoRA ($\tau=25$) 75.42 36.45 65.22 25.78 93.35 74.16 57.09 26.48 74.59 21.05 86.18 37.64 85.30 44.58 73.09 28.44 68.52 33.50
AdvCLIP-LoRA ($\tau=30$) 75.47 36.78 65.18 26.00 93.47 74.28 57.51 27.01 74.70 21.11 86.26 38.62 84.90 44.62 73.12 28.97 68.62 33.63
Relative Improvement  +$18.91$  +$94.87$  +$29.57$  +$39.09$  +$2.17$  +$30.73$  +$26.07$  +$252.02$  +$25.01$  +$85$  +$18.57$  +$143.64$  +$26.57$  +$62.09$  +$12.82$  +$137.71$  +$15.69$  +$67.05$
1 AdvVP Mao et al. 43.62 17.94 46.60 11.07 85.73 50.33 26.97 12.93 24.43 5.23 57.60 22.73 63.10 29.70 3.37 0.40 41.20 11.10
APT Li et al. (2024) 59.07 5.08 49.30 1.30 84.77 26.90 41.67 3.83 56.57 0.83 70.23 0.60 61.97 2.10 54.50 3.87 53.53 1.23
AdvPT Zhang et al. (2024) 29.96 1.44 20.17 0.43 62.97 7.60 16.73 2.60 13.27 0.00 37.93 0.13 33.97 0.43 27.03 0.00 27.57 0.37
AdvMaPLe Khattak et al. (2023) 33.52 11.38 49.27 14.60 85.53 48.37 13.63 2.93 5.27 0.30 30.67 4.97 1.40 0.10 32.70 7.07 49.70 12.67
AdvVLP Zhou et al. (2024) 33.01 12.03 50.53 17.50 85.43 48.47 15.97 4.77 1.07 0.77 29.63 3.83 19.77 6.57 11.83 1.73 49.83 12.60
FAP Zhou et al. (2024) 40.14 10.22 49.90 15.40 83.53 41.13 18.40 2.40 31.67 1.43 49.23 3.47 10.40 0.53 28.50 2.43 49.53 14.93
AdvCLIP-LoRA ($\tau=1$) 73.87 29.83 65.17 20.48 93.10 66.00 50.24 18.38 78.91 15.72 87.03 35.57 76.25 34.35 72.27 21.73 68.00 26.45
AdvCLIP-LoRA ($\tau=2$) 73.80 30.30 65.28 20.89 93.06 66.49 49.65 18.68 78.90 16.41 86.97 36.03 76.17 34.55 72.48 22.34 67.92 26.99
AdvCLIP-LoRA ($\tau=5$) 73.91 31.50 65.13 22.20 93.27 67.30 50.12 19.21 78.89 17.39 87.14 36.93 76.09 36.87 72.59 23.47 68.06 28.65
AdvCLIP-LoRA ($\tau=10$) 73.75 32.25 64.96 23.02 93.51 68.19 48.82 19.03 78.81 18.41 87.14 37.59 76.21 37.64 72.40 24.45 68.13 29.68
AdvCLIP-LoRA ($\tau=15$) 73.80 32.99 64.88 23.69 93.47 69.17 49.53 19.44 78.73 19.07 87.33 38.46 76.05 38.37 72.40 25.17 68.03 30.57
AdvCLIP-LoRA ($\tau=20$) 73.83 33.55 64.68 24.04 93.39 69.61 50.12 19.98 78.72 19.64 87.00 39.33 76.37 38.69 72.24 25.75 68.11 31.38
AdvCLIP-LoRA ($\tau=25$) 73.66 33.88 64.65 24.29 93.23 69.98 49.88 20.04 78.65 20.09 87.14 39.30 75.64 39.18 71.82 26.38 68.25 31.80
AdvCLIP-LoRA ($\tau=30$) 73.60 34.28 64.60 24.52 93.31 70.34 49.76 20.80 78.58 20.26 86.84 40.15 75.68 39.38 71.95 26.78 68.10 32.01
Relative Improvement  +$24.85$  +$79.77$  +$28.56$  +$31.54$  +$9.08$  +$35.49$  +$17.16$  +$47.18$  +$39.31$  +$252.01$  +$24.08$  +$65.38$  +$20.78$  +$26.73$  +$32.84$  +$245.83$  +$27.27$  +$98.79$
This section expands Section 5.2 with full adversarial few-shot results on ViT-B/32 across $\{1,2,4,8,16\}$ shots. In addition to baseline comparisons, we sweep the number of inner maximization steps $\tau$ used to train AdvCLIP-LoRA. Baselines report training with a 2-step PGD procedure, and the impact of using more steps is unspecified; therefore, for fairness, the main paper fixes $\tau=2$. The extended tables here show that increasing $\tau$ consistently improves AdvCLIP-LoRA’s robustness and overall accuracy. Here, we also report $\Delta$, defined as the relative improvement of AdvCLIP-LoRA with $\tau=10$ over the strongest non-ours baseline for each dataset and shot.
B.2 Comparative Analysis of AdvCLIP-LoRA and CLIP-LoRA